if the mass of water filling bottle is 25g and the mass of oil filling the bottle is 20g. What is the density of the oil.


1
Expert's answer
2022-01-23T16:14:00-0500

Given:

"m_1=25\:\rm g,\quad \rho_1=1000\: kg/m^3"

"m_2=20\:\rm g,\quad \rho_2=?"


The mass is dependent on density and volume by en equation

"m=\rho V"

Two liquids are filled at same bottle (volume), so

"\frac{m_1}{\rho_1}=\frac{m_2}{\rho_2}"

"\rho_2=\frac{m_2}{m_1}\rho_1=\frac{20}{25}*1000\:\rm kg/m^3=800\: kg/m^3"